OpenStudy (anonymous):

Is 4y+8x=1 in standard form? If not, put it in standard form and show your steps

OpenStudy (anonymous):

set the equation to 0. So move the 1 to the left side.

OpenStudy (anonymous):

is it 4y+8x=1 and simplified down to y=1/4-2x

OpenStudy (anonymous):

nothing to do here.........*leaves*

OpenStudy (anonymous):

or y=-2x-1/4

OpenStudy (anonymous):

y = -2x -1/4 is the slope y intercept form. You want the standard form

OpenStudy (anonymous):

The standard form is like this: Ax + By + C = 0

OpenStudy (anonymous):

ok so this equation is in standard form if u put it as 8x+4y-1=0

OpenStudy (anonymous):

yes